# From Single Genes to Populations: Quantifying Broken Detailed Balance in Transcription

### v2

@@ Abstract
Are there thermodynamic or information theoretic constraints governing mRNA expression? Here we use the framework of stochastic thermodynamics to assess detailed balance breaking in transcription. For the canonical two-state model of transcription, we derive exact analytic expressions for the entropy production rate of transcription at steady state, expressions that can be evaluated from knowledge of the kinetic parameters of the two-state model. This allows us to easily evaluate the entropy production rate of thousands of genes across seven datasets of two-state model parameters without needing to evaluate the entropy production rate from trajectory-based computation. A data-driven approach then exposes that most genes avoid parameter regimes associated with large entropy production rates, akin to a mesoscopic version of energy expenditure minimization. Importantly, we show that this is not a thermodynamic phenomenon, since the entropy production rate from the two state gene model provides only a weak bound on the housekeeping energy needed to power transcription. Finally, we show that cell-to-cell variability can make mRNA expression seem more or less irreversible than a “representative cell” would imply. Overall, our study presents a theoretical and data-driven approach to uncovering the potential thermodynamic and information theoretic constraints that dictate observed behaviors in transcription.
@@
Cells are fundamentally non-equilibrium systems [1]. They metabolize free energy to drive processes such as transcription, translation, regulation, and replication—activities necessary for survival and reproduction. If free energy sources are not present, mRNA and protein production ceases. Eventually, the system relaxes toward equilibrium with its environment, and the cell ultimately dies. In a proliferating population, this non-equilibrium character is evident in the irreversibility of exponential growth [2, 3, 4]. Yet, the extent to which the molecular processes underpinning this irreversibility violate detailed balance—and the degree to which reversibility is a determining factor in the dynamics of molecular machinery—remains an open question [5].
@@
In this paper, we take interest in understanding how the behaviors exhibited in the transcription of genes may be related to the ways in which those behaviors break time-reversal symmetry—otherwise known as detailed balance. Fundamentally for our analyses, we model transcription as a noisy process that is far-from-equilibrium [6, 7, 8]. This noise arises not only from extrinsic factors that vary between cells but also from the random waiting times between molecular collisions among genes, mRNA, and protein molecules [9, 10, 11, 12, 13]. The most common models of transcription are phenomenological—namely, stochastic switching of the gene between active (on) and inactive (off) states, transcription occurring only when the gene is active, and the subsequent degradation of mRNA transcripts. Simple models are highly successful at capturing the dynamics of genes that have mRNA distributions that are noisier than a Poisson point process [14]. Under these conditions, gene expression is said to be bursty, being characterized by short on-times and long waits in between bursts of mRNA production. This “bursty” nature is studied both at the level of transcript distributions across population snapshots [15, 16, 17, 18, 19] and in time for individual genes [20, 21, 22, 23].
@@
Figure 1: Schematics of the models considered in this study. (a) Illustration of a eukaryotic cell with the most important molecules and processes labeled. The gray box shows the model of transcription considered in this study. (b) Typical modeling and inference schemes assume populations of identical cells, and that most of cellular noise is intrinsic. A population of heterogeneous cells including varying features such as cell size or activator concentrations, leading to distributions of kinetic parameters over the populations [24].
@@
Stochastic thermodynamic analyses aimed at quantifying the degree to which mRNA expression dynamics violate detailed balance are relatively uncommon [25, 26, 27, 28, 6]. In part, this is because the common models of transcription are phenomenological, and Markov state transitions do not correspond to individual energy barriers. For instance, in models of constitutive expression—where a gene is presumed to produce mRNA in a Poisson process—the underlying process is, in reality, composed of many elementary steps describing the stepping of RNA Pol II along the gene. The Poissonian steady state described by constitutive expression, i.e., $G\to G+M,\;M\to\varnothing$, although in detailed balance is not an equilibrium because the reservoirs containing the ATP to fuel the process are chemostatted [29]. One principled way to bridge this gap between macroscopic models and microscopic thermodynamics is to consider the entropy production of a coarse-grained model as a lower bound to the entropy produced by the fine-grained system [30, 31, 32, 2, 33, 34, 35]. However, such bounds may not always be tight [36], and in some cases can even be violated if care is not taken in how decay products and reactants are distinguished [37].
@@
Rather than asking what the thermodynamics of a coarse-grained process can reveal about a more fine-grained process, one can instead consider the reverse: how does mesoscopic or macroscopic irreversibility emerge from the irreversibility or more fine-grained behavior? For example, in [38, 39], Lynn et al. devise algorithms to explore how exposing human subjects to different tasks changes the level of broken detailed balance in the brain, making the case that violations of detailed balance are necessary for cognition. From this perspective, one can ask how the irreversibility of a proliferating population of cells may arise from the dynamics of individual genes, and how the reality of cell-to-cell variability may affect these interpretations. A major question this paper tackles is: Is dynamical irreversibility a critical feature of transcriptional regulation—one that is necessary for robust cellular function?
@@
In this work, we develop simple analytic tools to study the breaking of detailed balance in transcription, under the framework of a two-state model of gene expression. Our analytic expressions allow entropy production rates to be directly computed from a gene’s kinetic parameters, allowing us to avoid simulation-based methods of calculating dynamical reversibility at a trajectory level. We apply these tools to estimate entropy production across thousands of genes in mice, across seven datasets of inferred two-state model parameters. Although the thermodynamic bounds provided by the telegraph model’s entropy production rate are not tight, we show that real genes choose kinetic parameters that clearly avoid areas of high entropy production. Finally, we assess analytically how cell-to-cell variability affects the irreversibility of gene expression at the population level, and the potential implications this has in determining where the cell may tolerate cell-to-cell variability.
I Two-state model of transcription
@@
Although there is substantial evidence suggesting that transcriptional kinetics are more complex than a simple two-state Markovian gene model implies [21, 40, 41, 16, 42, 43, 44], all widely accessible parameter inference datasets have relied on a two-state framework. In this model, key parameters such as the gene’s activation rate, inactivation rate, and transcription rate have been estimated across thousands of genes [45, 46, 19]. Generally, the two-state model is considered under the further assumption that a gene admits no “leakiness”, and that a transcriptionally off-state cannot be transcribed—this model is known as the telegraph model [47, 8, 48]. Two-state gene models are similar to models in the physics literature that have studied the stochastic dynamics of systems with switching environments [49].
@@
A major advantage of the two-state model is its distinguishability: from population level mRNA sequencing of mRNA FISH experiments, its three kinetic parameters can be robustly inferred. In contrast, comprehensive databases of kinetic parameters inferred from models with more than two gene states do not yet exist, in part because of the difficulty in performing experiments that would allow for transitions between gene states to be distinguished on a the level of a single cell [16, 50, 51]. Other complications in parameter inference include the reality that transcriptional kinetics are also dependent on factors that vary from cell-to-cell, and it has been concluded by some that mRNA transcript variability is determined by the properties of individual cells and not noise intrinsic to transcription and translation [52, 53, 24].
@@
[…]
$G$ and $G^{\star}$ represent the two different gene states with $n_{G}+n_{G^{\star}}=1$ (there is a single gene) and $M$ represents mRNA. This is also known as the leaky telegraph model or leaky gene model which was first solved in the steady state regime in [54]. Standard narratives ascribe $\rho_{\mathrm{off}}=0$ such that $G$ is a true off state. However, this assumption is not required, and we arbitrarily set $\rho_{\mathrm{on}}>\rho_{\mathrm{off}}$. It is well established that gene repression is not perfect, leading to basal expression—across the tree of life [55, 56, 57]. In practice, two-state models coarse-grain over a wide range of underlying molecular processes, and in the general case, the off state may simply exhibit a reduced transcription rate rather than none at all. This leads to a mesoscopic model of transcriptional dynamics that is dynamically reversible—also known as weak reversibility—although the dynamics are generally not reversible in the thermodynamic sense. A schematic of this process is shown in Fig. 1(a). The turning on (off) of the gene state could be related to the binding (unbinding) of an abundant activating transcription factor (and vice versa for a repressor), and the loss of $M$ is due to degradation.
@@
Below, $\mathbf{x}=(n_{G},n)$ denotes the state vector describing the number of $G$ and $M$ respectively (with the number of $G^{\star}$ being $1-n_{G}$), this set of reactions has the Markov state diagram shown in Fig. 2. A steady state that admits detailed balance is only guaranteed for one-dimensional reversible Markov chains, or for dynamically reversible reactions in closed vessels with specific parameter choices (see the final equation in [58] and generally [59]). The probabilities $P_{0}(n)$ and $P_{1}(n)$ denoting the probabilities of having states $(0,n)$ and $(1,n)$ are described by the master equations,
@@
and $\mathbb{E}^{x}$ is the step operator that acts such that $\mathbb{E}^{x}f(n)=f(n+x)$ [59]. One can then introduce the generating functions $G_{0}(z)=\sum_{n}z^{n}P_{0}(n)$ and $G_{1}(z)=\sum_{n}z^{n}P_{1}(n)$ and solve the resulting ODE at steady state when $\partial_{t}G_{0}=\partial_{t}G_{1}=0$ to find,
@@
This solution has also been provided in [54], although $G_{0}(z)$ and $G_{1}(z)$ are not provided therein. For later use, one can show that in the “bursty limit” of $\{\rho_{\mathrm{on}},\sigma_{\mathrm{off}}\}\gg\{\sigma_{\mathrm{off}},\rho_{\mathrm{off}},d\}$ that the Fano factor ($\mathrm{FF}$) and coefficient of variation squared ($CV^{2}$) of mRNA molecule numbers are given by,
@@
Note that these expressions are also calculable in the non-bursty regime, but the expressions are more complicated. These expressions can be found in Sec. S1.
@@
To understand the non-equilibrium nature of the two-state gene, one can use the equations for $G_{0}(z)$ and $G_{1}(z)$ to calculate the probability fluxes between the Markov states in Fig. 2. First, it is clear that if $\rho_{\mathrm{off}}=\rho_{\mathrm{on}}$ then the gene states cannot be distinguished and the generating function reduces to that of a Poisson process, because ${}_{1}F_{1}(a,b;0)=1$. Typically, in the context of bursty gene expression (where $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$) $\sigma_{\mathrm{on}}$ is known as the burst frequency and $B=\rho_{\mathrm{on}}/\sigma_{\mathrm{off}}$ as the mean burst size. Although there is no commonly accepted definition of how bursty a gene is, here we use two measures: (i) $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$, which is the ratio of timescale of off switching to on switching—larger $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$ implies a greater degree of burstiness because more time is spent in a transcriptionally impaired state; (ii) $CV^{2}$, because bursting is often thought of as confining variability to the level of mRNA transcripts, and $CV^{2}$ is a direct measure of transcript variability.
@@
Figure 3: Exploring the non-equilibrium steady state of the two-state model. (a) Plots of $P_{0}(n)$, $P_{1}(n)$ and $P(n)$ calculated using the generating functions in the main text for parameters $\rho_{\mathrm{off}}=1/3,\rho_{\mathrm{on}}=30,\sigma_{\mathrm{off}}=\sigma_{\mathrm{on}}=1/2,d=1$. (b) A diagram showing the flow of probability flux between states in the Markov diagram. Variables $b_{n}$ and $a_{n}$ represent probability fluxes, whose expressions are given in the main text. Here $b_{n}>0$, whereas the $a_{n}$ can be positive or negative.
@@
Fig. 3 shows a probability flux analysis on the two-state model. To get some intuition for the nature of the steady state one can look at the net probability flux between each pair of connected states, shown in Fig. 3(b) [60]. In words, there is a unidirectional flux of probability from right-to-left along the states $(0,n)\to(0,n-1)\,\forall\,n\in\{1,2,\ldots\}$ and from left-to-right along the states $(1,n-1)\to(1,n)\,\forall\,n\in\{1,2,\ldots\}$ which in both cases is of magnitude $b_{n}>0$. Between the two chains there are fluxes $a_{n}$ between states $(0,n)$ and $(1,n)$ which can be positive or negative, with the exception of $a_{0}=b_{1}>0$ to conserve probability at the end of the chain.
@@
An interesting observation is that if one ignores the state of the gene, on the level of the mRNA detailed balance is maintained. This is also true if one were to lump together the mRNA states and only observe the gene state dynamics, since one can show that $-b_{1}=\sum_{i=1}^{\infty}a_{i}$. This finding, that lumping gene states gives marginal dynamics satisfying detailed balance, aligns with other recent findings that non-equilibrium systems with hidden states can “pretend” to obey detailed balance [35]. Marginal detailed balance on the mRNA number means that the condition,
@@
must be true, which can be shown directly from the generating functions above. This marginal detailed balance condition is hinted to by the fact that the eigenvalues governing the relaxation to the steady state distribution are real [59, 61], given by the two sets $\lambda^{1}_{n}=nd,\;n\in\{0,1,2,\ldots\}$ and $\lambda^{2}_{n}=nd+\Sigma,\;n\in\{0,1,2,\ldots\}$, which is shown in Section S2. Note that, although it is true that a system satisfying detailed balance will have a master operator with real eigenvalues, the converse is generally not true. This two-state gene model is an example of this fact. We explore further implications of marginal detailed balance in Section S5.
II Entropy production of transcription in a two-state gene
To investigate the degree to which an individual gene—comprising kinetic parameters $\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},\rho_{\mathrm{on}},\rho_{\mathrm{off}}$ and $d$—breaks detailed balance, we can utilize a quantity known as the entropy production rate (EPR) [62, 63, 28]. The EPR has a couple of very useful properties. First, the EPR tells us the degree to which a process is irreversible. For a system in detailed balance, the EPR is zero because every forward transition has an equal reverse probability flux—this is the requirement for time-reversal symmetry. Zero is the minimum value of the EPR. For a dynamically irreversible system, the EPR is infinite because not every forward transition has a reverse transition—this is the requirement to have forward trajectories that could never be seen in reverse. Second, for a microscopic system made up of elementary reactions, the EPR is proportional to the amount of free energy needed to power the system in its non-equilibrium dynamics [64, 65]. The aims of this section are: (i) to find an analytic formula for the EPR of a transcribing gene; (ii) to apply this formula across several kinetic parameter datasets for genes in mouse cells; (iii) to understand the degree to which this EPR gives a thermodynamic lower bound on the microscopic EPR; and (iv) to discover, from a data-driven approach, whether reversibility is a desirable property in the dynamics of mRNA expression.
Following standard methods [62, 28], the EPR is given by (in natural units),
@@ II Entropy production of transcription in a two-state gene
The EPR in Eq. (7) is often broken down into components of entropy flow rate—entropy production due to interactions between the system and an external reservoir of energy or particles— and the entropy change of the system—relating to changes in the internal configurations of the system [60]. However, we assume the system is at a steady state, meaning that the EPR is simply the negative of the entropy flow rate. It is equal to the negative entropy flow rate due to the convention in stochastic thermodynamics that it is the system doing the work, and not having work done on it. Negative entropy flow therefore implies that the system is subject to influx from the reservoirs. The equation for the entropy flow rate, $\dot{s}_{e}$, is (in natural units),
@@
For further details on the relationship between the entropy production rate and the entropy flow rate at steady state, please refer to Section S3, and references [60, 27, 66, 67].
@@
Eq. (10) elucidates that if either $\sigma_{\mathrm{on}}$ or $\sigma_{\mathrm{off}}$ are zero then the steady state is also in detailed balance since one is coupled to only a single mRNA reservoir. This formula also tells us that as $\rho_{\mathrm{off}}\to 0$ the entropy production becomes infinite, and the process becomes completely irreversible—albeit logarithmically slowly. Therefore, the often employed assumption of $\rho_{\mathrm{off}}=0$ actually represents the most non-equilibrium form of the two-state model [45, 21], as is clear from Crooks’ fluctuation theorem [68]. Curiously, when $\rho_{\mathrm{off}}=0$, even though the dynamics are completely irreversible, the marginal mRNA dynamics still satisfy detailed balance at steady state. This fact should provide a motivation for the study of coarse graining in mechanistic mRNA expression models that have an effective off state without being thermodynamically impossible—as coarse graining via lumping should give effective models that provide a lower bound on the EPR, not an infinite value of the EPR [69].
@@
  • Two datasets (both of around 1,500 genes) split into G1 and G2M cell-cycle stages in mouse fibroblasts (from [19], based on data in [70]).
@@
  • One recent dataset (of around 5,600 gene) in mouse fibroblasts [46].
@@
  • Four datasets (each of around 5,000-6,000 genes) across both mouse fibroblasts and mouse embryonic stem cells [45].

### v3

# Quantifying Broken Detailed Balance in Transcription

@@ Abstract
For the canonical two-state model of transcription, we derive exact analytic expressions for the entropy production rate of transcription at steady state, and assess detailed balance breaking in transcription. Our analytics allow us to easily evaluate the entropy production rate of thousands of genes across seven datasets of two-state model parameters without needing to evaluate the entropy production rate from trajectory-based computation. A data-driven approach then exposes that most genes avoid parameter regimes associated with large entropy production rates, akin to a mesoscopic version of energy expenditure minimization. Importantly, we show that this is not a thermodynamic phenomenon, since the entropy production rate from the two state gene model provides only a weak bound on the housekeeping energy needed to power transcription. Finally, we show that cell-to-cell variability can make mRNA expression seem more or less irreversible than a “representative cell” would imply.
Introduction
@@
Cells are fundamentally non-equilibrium systems [schrodinger2012life]. They metabolize free energy to drive processes such as transcription, translation, regulation, and replication—activities necessary for survival and reproduction. If free energy sources are not present, mRNA and protein production ceases. Eventually, the system relaxes toward equilibrium with its environment, and the cell ultimately dies. In a proliferating population, this non-equilibrium character is evident in the irreversibility of exponential growth [england2013statistical, grandpre2025extremal, ocal2025two]. Yet, the extent to which the molecular processes underpinning this irreversibility violate detailed balance—and the degree to which reversibility is a determining factor in the dynamics of molecular machinery—remains an open question [battle2016broken].
@@
In this paper, we take interest in understanding how the behaviors exhibited in the transcription of genes may be related to the ways in which those behaviors break time-reversal symmetry—otherwise known as detailed balance. Fundamentally, transcription is a noisy process that is far-from-equilibrium [zoller2022eukaryotic, wong2020gene, raj2006stochastic]. This noise arises not only from extrinsic factors that vary between cells but also from the random waiting times between molecular collisions among genes, mRNA, and protein molecules [mcadams1997stochastic, elowitz2002stochastic, swain2002intrinsic, raj2008nature, VERHAGEN2025169202]. The most common models of transcription are phenomenological—namely, stochastic switching of the gene between active (on) and inactive (off) states, transcription occurring only when the gene is active, and the subsequent degradation of mRNA transcripts. Simple models are highly successful at capturing the dynamics of genes that have mRNA distributions that are noisier than a Poisson point process [braichenko2021distinguishing]. Under these conditions, gene expression is said to be bursty, being characterized by short on-times and long waits in between bursts of mRNA production. This “bursty” nature is studied both at the level of transcript distributions across population snapshots [molina2013stimulus, bartman2019transcriptional, volteras2023global, fu2022quantifying, sukys2025cell] and in time for individual genes [golding2005real, suter2011mammalian, lammers2020matter, pomp2024transcription].
@@
Figure 1: Schematics of the models considered in this study. (a) Illustration of a eukaryotic cell with the most important molecules and processes labeled. The gray box shows the model of transcription considered in this study. (b) Typical modeling and inference schemes assume populations of identical cells, and that most of cellular noise is intrinsic. A population of heterogeneous cells including varying features such as cell size or activator concentrations, leading to distributions of kinetic parameters over the populations [lim2015quantitative, grima2023quantifying].
@@
Stochastic thermodynamic analyses aimed at quantifying the degree to which mRNA expression dynamics violate detailed balance are relatively uncommon [tkavcik2008information, gama2020binary, boeger2022energetics, boeger2022kinetic, zoller2022eukaryotic]. In part, this is because the common models of transcription are phenomenological, and Markov state transitions do not correspond to individual energy barriers. For instance, in models of constitutive expression—where a gene is presumed to produce mRNA in a Poisson process—the underlying process is, in reality, composed of many elementary steps describing the stepping of RNA Pol II along a gene. The Poissonian steady state described by constitutive expression, i.e., $G\to G+M,\;M\to\varnothing$, although in detailed balance is not an equilibrium because the reservoirs containing the ATP to fuel the process are chemostatted [qian2007phosphorylation]. One principled way to bridge this gap between macroscopic models and microscopic thermodynamics is to consider the entropy production of a coarse-grained model as a lower bound to the entropy produced by the fine-grained system [egolf2000equilibrium, blythe2008reversibility, gomez2008lower, england2013statistical, bo2017multiple, jia2016model, skinner2021estimating]. However, such bounds may not always be tight [ghosal2022inferring], and in some cases can even be violated if care is not taken in how decay products and reactants are distinguished [kolchinsky2024thermodynamic].
@@
Rather than asking what the thermodynamics of a coarse-grained process can reveal about a more fine-grained process, one can instead consider the reverse: how does mesoscopic or macroscopic irreversibility emerge from the irreversibility or more fine-grained behavior? For example, in [lynn2021broken, lynn2022emergence], Lynn et al. devise algorithms to explore how exposing human subjects to different tasks changes the level of broken detailed balance in the brain, making the case that violations of detailed balance are necessary for cognition. From this perspective, one can ask how the irreversibility of a proliferating population of cells may arise from the dynamics of individual genes, and how the reality of cell-to-cell variability may affect these interpretations. A major question this paper tackles is: how does the irreversible nature of cellular dynamics propagate down to the level of individual genes?
@@
In this work, we develop simple analytic tools to study the breaking of detailed balance in transcription, under the framework of a two-state model of gene expression. Our analytic expressions allow entropy production rates to be directly computed from a gene’s kinetic parameters, allowing us to avoid simulation-based methods of calculating dynamical reversibility at a trajectory level. We apply these tools to estimate entropy production across thousands of genes in mice, across seven datasets of inferred two-state model parameters. Although the thermodynamic bounds provided by the telegraph model’s entropy production rate are not tight, we show that real genes choose kinetic parameters that clearly avoid areas of high entropy production. Defining two statistical null models we show that similar properties are not as pronounced when transcriptional kinetic parameters are randomized. Finally, we assess analytically how cell-to-cell variability affects the irreversibility of gene expression at the population level, and the potential implications this has in determining where the cell may tolerate cell-to-cell variability.
Two-state model of transcription
@@
Although there is substantial evidence suggesting that transcriptional kinetics are more complex than a simple two-state Markovian gene model implies [suter2011mammalian, bothma2014dynamic, rodriguez2019intrinsic, bartman2019transcriptional, cao2020stochastic, tunnacliffe2020transcriptional, shelansky2024single], all widely accessible parameter inference datasets have relied on a two-state framework. In this model, key parameters such as the gene’s activation rate, inactivation rate, and transcription rate have been estimated across thousands of genes [larsson2019genomic, ramskold2024single, sukys2025cell]. Generally, the two-state model is considered under the further assumption that a gene admits no “leakiness”, and that a transcriptionally off-state cannot be transcribed—this model is known as the telegraph model [peccoud1995markovian, raj2006stochastic, iyer2009stochasticity]. Two-state gene models are similar to models in the physics literature that have studied the stochastic dynamics of systems with switching environments [hufton2016intrinsic].
@@
A major advantage of the two-state model is its distinguishability: from population level mRNA sequencing of mRNA FISH experiments, its three kinetic parameters can be robustly inferred. In contrast, comprehensive databases of kinetic parameters inferred from models with more than two gene states do not yet exist, in part because of the difficulty in performing experiments that would allow for transitions between gene states to be distinguished on a the level of a single cell [bartman2019transcriptional, wildner2023bayesian, eck2024single]. Other complications in parameter inference include the reality that transcriptional kinetics are dependent on factors that vary from cell-to-cell, and it has been concluded by some that mRNA transcript variability is determined by the properties of individual cells and not noise intrinsic to transcription and translation [battich2015control, ham2020extrinsic, grima2023quantifying].
@@
$G\xrightarrow{\rho_{\mathrm{off}}}G+M,\;G^{\star}\xrightarrow{\rho_{\mathrm{on}}}G^{\star}+M,\;G\xrightleftharpoons[\sigma_{\mathrm{off}}]{\sigma_{\mathrm{on}}}G^{\star},\;M\xrightarrow{d}\varnothing.$ (1)
$G$ and $G^{\star}$ represent the two different gene states with $n_{G}+n_{G^{\star}}=1$ (there is a single gene) and $M$ represents mRNA. This is also known as the leaky telegraph model or leaky gene model which was first solved in the steady state regime in [ham2020exactly]. Standard narratives ascribe $\rho_{\mathrm{off}}=0$ such that $G$ is a true off state. However, this assumption is not required, and we arbitrarily set $\rho_{\mathrm{on}}>\rho_{\mathrm{off}}$. It is well established that gene repression is not perfect, leading to basal expression—across the tree of life [phan2015development, zacharioudakis2016novel, flouriot2020basal]. In practice, two-state models coarse-grain over a wide range of underlying molecular processes, and in the general case, the off state may simply exhibit a reduced transcription rate rather than none at all. This leads to a mesoscopic model of transcriptional dynamics that is dynamically reversible—also known as weak reversibility—although the dynamics are generally not reversible in the thermodynamic sense. A schematic of this process is shown in Fig. 1(a). The turning on (off) of the gene state could be related to the binding (unbinding) of an abundant activating transcription factor (and vice versa for a repressor), and the loss of $M$ is due to degradation.
@@
Below, $\mathbf{x}=(n_{G}^{\star},n)$ denotes the state vector describing the number of $G^{\star}$ and $M$ respectively (with the number of $G^{\star}$ being $1-n_{G}$), this set of reactions has the Markov state diagram shown in Fig. 2. A steady state that admits detailed balance is only guaranteed for one-dimensional reversible Markov chains, or for dynamically reversible reactions in closed vessels with specific parameter choices (see the final equation in [van1976equilibrium] and generally [van1992stochastic]). The probabilities $P_{0}(n)$ and $P_{1}(n)$ denoting the probabilities of having states $(0,n)$ and $(1,n)$ are described by the master equations,
@@
and $\mathbb{E}^{x}$ is the step operator that acts such that $\mathbb{E}^{x}f(n)=f(n+x)$ [van1992stochastic]. One can then introduce the generating functions $G_{0}(z)=\sum_{n}z^{n}P_{0}(n)$ and $G_{1}(z)=\sum_{n}z^{n}P_{1}(n)$ and solve the resulting ODE at steady state when $\partial_{t}G_{0}=\partial_{t}G_{1}=0$ to find,
@@
This solution has also been provided in [ham2020exactly], although $G_{0}(z)$ and $G_{1}(z)$ are not provided therein. For later use, one can show that in the “bursty limit” of $\{\rho_{\mathrm{on}},\sigma_{\mathrm{off}}\}\gg\{\sigma_{\mathrm{off}},\rho_{\mathrm{off}},d\}$ that the Fano factor ($\mathrm{FF}$) and coefficient of variation squared ($CV^{2}$) of mRNA molecule numbers are given by,
@@
Note that these expressions are also calculable in the non-bursty regime, but the expressions are more complicated. Derivations of these expressions can be found in Sec. S1.
@@
To understand the non-equilibrium nature of the two-state gene, one can use the equations for $G_{0}(z)$ and $G_{1}(z)$ to calculate the probability fluxes between the Markov states in Fig. 2. First, it is clear that if $\rho_{\mathrm{off}}=\rho_{\mathrm{on}}$ then the gene states cannot be distinguished and the generating function reduces to that of a Poisson process, since $_{1}F_{1}(a,b;0)=1$. Typically, in the context of bursty gene expression (where $\rho_{\mathrm{off}}\ll\rho_{\mathrm{on}}$) $\sigma_{\mathrm{on}}$ is known as the burst frequency and $B=\rho_{\mathrm{on}}/\sigma_{\mathrm{off}}$ as the mean burst size. Although there is no commonly accepted definition of how bursty a gene is, here we use two measures: (i) $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$, which is the ratio of timescale of off switching to on switching—larger $\sigma_{\mathrm{off}}/\sigma_{\mathrm{on}}$ implies a greater degree of burstiness because more time is spent in a transcriptionally impaired state; and (ii) $CV^{2}$, because bursting is often thought of as confining variability to the level of mRNA transcripts, and $CV^{2}$ is a direct measure of transcript variability.
@@
Figure 3: Exploring the non-equilibrium steady state of the two-state model. (a) Plots of $P_{0}(n)$, $P_{1}(n)$ and $P(n)$ calculated using the generating functions in the main text for parameters $\rho_{\mathrm{off}}=1/3,\rho_{\mathrm{on}}=30,\sigma_{\mathrm{off}}=\sigma_{\mathrm{on}}=1/2,$ and $d=1$. (b) A diagram showing the flow of probability flux between states in the Markov diagram. Variables $b_{n}$ and $a_{n}$ represent probability fluxes, whose expressions are given in the main text. Here $b_{n}>0$, whereas the $a_{n}$ can be positive or negative.
@@
Fig. 3 shows a probability flux analysis on the two-state gene model. To get some intuition for the nature of the steady state one can look at the net probability flux between each pair of connected states, shown in Fig. 3(b) [schnakenberg1976network]. In words, there is a unidirectional flux of probability from right-to-left along the states $(0,n)\to(0,n-1)\,\forall\,n\in\{1,2,\ldots\}$ and from left-to-right along the states $(1,n-1)\to(1,n)\,\forall\,n\in\{1,2,\ldots\}$ which in both cases is of magnitude $b_{n}>0$. Between the two chains there are fluxes $a_{n}$ between states $(0,n)$ and $(1,n)$ which can be positive or negative, with the exception of $a_{0}=b_{1}>0$ to conserve probability at the end of the chain.
@@
An interesting observation is that if one ignores the state of the gene, on the level of the mRNA detailed balance is maintained. This is also true if one were to lump together the mRNA states and only observe the gene state dynamics, since one can show that $b_{1}=\sum_{i=1}^{\infty}a_{i}$. This finding, that lumping gene states gives marginal dynamics satisfying detailed balance, aligns with other recent findings that non-equilibrium systems with hidden states can “pretend” to obey detailed balance [skinner2021estimating]. Marginal detailed balance on the mRNA number means that the condition,
@@
must be true, which can be shown directly from the generating functions above. This marginal detailed balance condition is hinted to by the fact that the eigenvalues governing the relaxation to the steady state distribution are real [van1992stochastic, tauber2014critical], given by the two sets $\lambda^{1}_{n}=nd,\;n\in\{0,1,2,\ldots\}$ and $\lambda^{2}_{n}=nd+\sigma_{\mathrm{on}}+\sigma_{\mathrm{off}},\;n\in\{0,1,2,\ldots\}$, which is shown in Section S2. Note that, although it is true that a system satisfying detailed balance will have a master operator with real eigenvalues, the converse is generally not true. This two-state gene model is an example of this fact. We explore further implications of marginal detailed balance in Section S3.
Entropy production of transcription in a two-state gene
To investigate the degree to which an individual gene—comprising kinetic parameters $\sigma_{\mathrm{on}},\sigma_{\mathrm{off}},\rho_{\mathrm{on}},\rho_{\mathrm{off}}$ and dd—breaks detailed balance, we can utilize a quantity known as the entropy production rate (EPR) [peliti2021stochastic, seifert2012stochastic, boeger2022kinetic]. The EPR has a couple of very useful properties. First, the EPR tells us the degree to which a process is irreversible. For a system in detailed balance, the EPR is zero because every forward transition has an equal reverse probability flux—this is the requirement for time-reversal symmetry. Zero is the minimum value of the EPR. For a dynamically irreversible system, the EPR is infinite because not every forward transition has a reverse transition—this is the requirement to have forward trajectories that could never be seen in reverse. Second, for a microscopic system made up of elementary reactions, the EPR is proportional to the amount of free energy needed to power the system in its non-equilibrium dynamics [ge2013dissipation, yoshimura2023housekeeping]. The aims of this section are: (i) to find an analytic formula for the EPR of a transcribing gene; (ii) to apply this formula across several kinetic parameter datasets for genes in mouse cells; (iii) to understand the degree to which this EPR gives a thermodynamic lower bound on the microscopic EPR; and (iv) to discover, from a data-driven approach, whether reversibility is a desirable property in the dynamics of mRNA expression.
Following standard methods [peliti2021stochastic, boeger2022kinetic], the EPR is given by (in natural units),
@@
The EPR in Eq. (7) is often broken down into components of entropy flow rate—entropy production due to interactions between the system and an external reservoir of energy or particles— and the entropy change of the system—relating to changes in the internal configurations of the system [schnakenberg1976network]. However, we assume the system is at a steady state, meaning that the EPR is simply the negative of the entropy flow rate. It is equal to the negative entropy flow rate due to the convention in stochastic thermodynamics that it is the system doing the work, and not having work done on it [peliti2021stochastic]. Negative entropy flow therefore implies that the system is subject to influx from the reservoirs. The equation for the entropy flow rate, $\dot{s}_{e}$, is (in natural units),
@@
For further details on the relationship between the entropy production rate and the entropy flow rate at steady state, please refer to Section S4, and references [schnakenberg1976network, boeger2022energetics, kirchberg2023energy, esposito2010three].
@@
Eq. (10) elucidates that if either $\sigma_{\mathrm{on}}$ or $\sigma_{\mathrm{off}}$ are zero then the steady state is also in detailed balance since one is coupled to only a single mRNA reservoir. This formula also tells us that as $\rho_{\mathrm{off}}\to 0$ the entropy production becomes infinite, and the process becomes completely irreversible—albeit logarithmically slowly. Therefore, the often employed assumption of $\rho_{\mathrm{off}}=0$ actually represents the most non-equilibrium form of the two-state model [larsson2019genomic, suter2011mammalian], as is clear from Crooks’ fluctuation theorem [crooks1999entropy]. Curiously, when $\rho_{\mathrm{off}}=0$, even though the dynamics are completely irreversible, the marginal mRNA dynamics still satisfy detailed balance at steady state. This fact should provide a motivation for the study of coarse graining in mechanistic mRNA expression models that have an effective off state without being thermodynamically impossible—as coarse graining via lumping should give effective models that provide a lower bound on the EPR, not an infinite value of the EPR [bo2014entropy].
@@
  • Two datasets (both of around 1,500 genes) split into G1 and G2M cell-cycle stages in mouse fibroblasts (from [sukys2025cell], based on data in [riba2022cell]).
@@
  • One recent dataset (of around 5,600 genes) in mouse fibroblasts [ramskold2024single].
@@
  • Four datasets (each of around 5,000-6,000 genes) across both mouse fibroblasts and mouse embryonic stem cells [larsson2019genomic].